# An FPGA-Based Open-Source Hardware-Software Framework for Side-Channel Security Research

### v1

@@ Abstract
Attacks based on side-channel analysis (SCA) pose a severe security threat to modern computing platforms, further exacerbated on IoT devices by their pervasiveness and handling of private and critical data. Designing SCA-resistant computing platforms requires a significant additional effort in the early stages of the IoT devices’ life cycle, which is severely constrained by strict time-to-market deadlines and tight budgets. This manuscript introduces a hardware-software framework meant for SCA research on FPGA targets. It delivers an IoT-class system-on-chip (SoC) that includes a RISC-V CPU, provides observability and controllability through an ad-hoc debug infrastructure to facilitate SCA attacks and evaluate the platform’s security, and streamlines the deployment of SCA countermeasures through dedicated hardware and software features such as a DFS actuator and FreeRTOS support. The open-source release of the framework includes the SoC, the scripts to configure the computing platform, compile a target application, and assess the SCA security, as well as a suite of state-of-the-art SCA attacks and countermeasures. The goal is to foster its adoption and novel developments in the field, empowering designers and researchers to focus on studying SCA countermeasures and attacks while relying on a sound and stable hardware-software platform as the foundation for their research.
@@ I Introduction
Hundreds of billions of Internet of Things (IoT) devices are getting more and more able to autonomously make decisions thanks to artificial intelligence (AI). As they get more capable, pervasive, connected, and able to interact with each other through 5G and 6G networks, IoT devices are constantly collecting, processing, and exchanging personal, private, and critical data, making it paramount to consider security as an essential aspect during their design and whole life cycle [1].
Relying on the theoretical security provided by traditional cryptography solutions such as AES, RSA, and SHA-3 and by secure communication protocols such as TLS and SSH is, however, insufficient to guarantee IoT device security. Side-channel analysis (SCA) attacks, in particular, target information collected from executing a specific implementation of a cryptographic scheme or security protocol rather than flaws in their specification or the theoretical problems at their foundation [2]. Such so-called side-channel information ranges from the time required for a cryptography-related computation to the power consumption, electromagnetic emissions, and noise produced during its execution. Moreover, while some fault-injection-based attacks interfere with the correct functioning of the target device, many other so-called passive side-channel attacks, such as differential power analysis (DPA) ones, are undetectable by the system under attack.
While SCA attacks increasingly become serious security threats to IoT devices, and in particular to those that operate in public spaces and are accessible by anyone, the computing platforms that power these devices and that handle personal, private, and critical data are, however, often not designed to protect from them. The commercially available microcontroller-based system-on-chips (SoCs) that provide the computing and connection capabilities of the IoT devices indeed commonly execute cryptography applications, while not being meant to do so in a secure way, mostly due to cost-saving reasons [3].
At the same time, the strict time-to-market deadlines, tight energy and power consumption and area constraints, and rapidly evolving tasks to be executed on such SoCs make it harder to offload their tasks, including cryptographic ones, to dedicated hardware accelerators and favor instead the design of simpler and more flexible programmable architectures that can also run an operating system (OS).
Conversely, the open literature does not provide a comprehensive hardware-software solution to tackle the challenge of SCA security on modern RISC-V platforms. The few RISC-V security-oriented solutions, e.g., OpenTitan [4], are not meant as general-purpose computing platforms and do not support the development of novel SCA attack and defense methodologies, while others focus their research effort solely on the hardware side while disregarding the software framework [5] or viceversa.
@@
This manuscript introduces an open-source research framework for side-channel analysis on FPGA-based IoT/microcontroller-class computing platforms, that encompasses a complete SoC and a software toolchain for SCA attacks and countermeasures, delivering a comprehensive hardware-software solution that is, to the best of our knowledge, currently lacking in the open literature.
The framework provides an IoT-class SoC that includes a CPU based on open-standard RISC-V as its instruction set architecture (ISA), dedicated hardware to enable the implementation of state-of-the-art SCA countermeasures, an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, and support for the open source FreeRTOS real-time operating system (RTOS). A complete flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify SCA vulnerabilities and leakage sources.
The goal of empowering designers and researchers to only focusing on studying SCA countermeasures is achieved through three main contributions:
  • Capability to identify leakage sources – Hardware security requires not just identifying the presence of a vulnerability, but also pinpointing the source of such leakage. The computing platform has a minimal architecture to expose the least leakage sources and make them eventually simpler to identify, in addition to being easier to emulate and cheaper in terms of area and power cost, while a one-to-one match between the prototyped platform and its emulated counterpart is enforced through dedicated hardware mechanisms to enable the collection of the most accurate side-channel signal.
  • Usability for research purposes – The framework is released open source, including all the hardware and software, from the SoC to the software scripts driving the SoC configuration, compilation, prototype execution and emulation, oscilloscope measurement, and state-of-the-art SCA attacks, to foster its adoption in research settings. Carrying out an analysis through the proposed framework provides not only the traditional power, performance, and area (PPA) quality metrics, but it extends them by adding the security dimension.
  • Complete SCA ecosystem – The framework includes dedicated hardware and software support for state-of-the-art SCA attacks and countermeasures that can be employed out of the box to directly compare with solutions from the literature as well as evaluate and enforce the SCA security of the computing platform. The experimental evaluation in this manuscript demonstrates the framework’s capabilities and showcases the SCA attacks and countermeasures included as part of its open-source release.
@@
The rest of this manuscript is organized into five parts. Section II provides a background on SCA attacks and countermeasures and an overview of the state-of-the-art RISC-V computing platforms. Section III describes the proposed framework and the hardware-software infrastructure that implements it. Section IV details the key hardware and software aspects that enable the SCA attack and countermeasure capabilities of the framework, Section V showcases the framework capabilities through a comprehensive experimental evaluation. Finally, Section VI draws conclusions and discusses the future works and developments on the proposed framework.
@@ II-A Side-channel analysis attacks and countermeasures
SCA attacks can be classified as either non-profiled or profiled. Non-profiled ones, such as those based on differential power analysis (DPA) [6] and correlation power analysis (CPA) [7], only target side-channel information that is obtained from the specific device under attack while it is performing a computation, and then attempt to recover the corresponding secret key by leveraging the statistical correlation between the measured side-channel signal and the data being processed, exploiting a partial knowledge of the latter. Profiled attacks [8] are carried out instead by initially making use of a replica of the device to be attacked to identify and fine-tune the side-channel leakage model and later employing such model to attack the actual target device, under the assumption that the latter and its replica share identical or at least similar leakage models [9]. Whereas SCA attacks have traditionally employed statistical techniques, machine- and deep-learning approaches recently emerged as promising research avenues for more capable attacks, ranging from the identification of cryptographic operations from side-channel power traces [10] to learning-based non-profiled [11] and profiled [12] attacks.
@@
Figure 1: High-level flow of the proposed framework.
@@ II-B State-of-the-art RISC-V computing platforms
On the one hand, a variety of RISC-V-based microcontroller-class SoCs is available, such as PULPino [18] and GAP-8 [19], while ultra-low-power platforms also leveraged RISC-V in multi-core heterogeneous architectures, e.g., Mr. Wolf [20] and HERO [21]. Such solutions focus on delivering good performance at a low energy and power consumption, but do not cover security and in particular SCA-related aspects.
On the other hand, frameworks such as ESP [22] and Chipyard [23] have emerged in the literature for the agile development of heterogeneous multi- and many-core SoCs that feature both general-purpose CPU cores and hardware accelerators [24]. However, they focus on the modularity and composability of the obtained computing platforms and on the fast prototyping pf the latter on FPGA targets, while they do not consider instead any security aspects.
@@
Finally, the OpenTitan project [4] delivers a hardware root of trust (RoT) centered around a RISC-V core, not providing instead neither a complete SoC platform nor a hardware-software framework for SCA-related research.
There also exists a variety of both RISC-V resource-constrained and application-class cores in the open literature, including CORE-V CV32E40P [25] and CVA6 [26], UCB’s Rocket [27], and Frontgrade Gaisler NOEL-V [28], that are however of limited use, when considered on their own, in providing a full platform for SCA research. State-of-the-art architectural simulators such as gem5 [29] and RISC-V-specific GVSoC [30] operate instead at a high level of abstraction, drastically shrinking the time required to simulate a computing platform but conversely making them unsuitable to produce information that can be effectively exploited by SCA attacks.
@@ III Framework
Figure 2: Hardware-software infrastructure that implements the proposed framework.
The proposed framework enables thoroughly evaluating the vulnerability to SCA attacks of cryptographic applications executed on an IoT-class SoC that implements countermeasures through dedicated hardware support. The possibility to accurately match the emulated and prototype execution of the computing platform allows not only detecting information leakage, but even more importantly to identify its sources, which is paramount to fixing the vulnerabilities of the SoC during the design phase. Section III-A outlines the high-level flow for the proposed framework, while Section III-B discusses how such flow is implemented in a hardware-software infrastructure.
@@ III-A High-level flow
The proposed framework, depicted in Figure 1, can be described as a sequence of three main phases devoted to 1) configuringthe hardware-software setup, 2) measuringSCA-related information from simulation and prototype execution, and 3) analyzingsuch information to detect leakage and identify its sources. Remarkably, the open-source nature of the framework, as well as the adoption of standard languages for all the inputs and its internals, enable the user to tinker, adapt, and tailor it to their needs. Software scripts automate the execution of each of the three phases, which are discussed in more detail in the rest of this section.
@@ III-A1 Configure
The high-level flow starts with the configuration of the hardware and software parts of the target to be simulated and prototyped. The SoC is thus configured by the user and implemented according to the standard hardware design flow, while the applications and, optionally, the RTOS are compiled to produce an executable to be run by the CPU.
@@
During this phase, the source code for the applications provided by the user of the proposed framework is compiled, through a compiler toolchain for RISC-V, to be executed on the target platform, both in simulation and on the prototype. The compilation produces an executable file for the compiled applications and, optionally, the RTOS, meant to be executed by the computing platform both in simulation and on the prototype FPGA. The boundaries for the time window of interest in the execution of the application, e.g., the computation of a specific cryptographic kernel of which to evaluate SCA resistance, are also extracted from the executable file.
@@ III-A3 Analyze
The final phase of the proposed framework foresees the analysis of the measurements from the previous phase, i.e., the switching activity of the simulation and the power consumption trace of the board execution, both corresponding to the time window of interest, in order to detect whether there was any cryptographic information leakage as well as to identify the eventual sources of such leakage. A set of state-of-the-art SCA techniques, ranging from CPA and template to ML-based ones, is therefore provided to evaluate the SCA security of the platform and the application executed on top of it. Remarkably, the analysis phase also produces the traditional power, performance, and area (PPA) metrics.
@@ III-B Hardware-software infrastructure
The hardware-software infrastructure depicted in Figure 2 realizes the high-level flow described in Section III-A by means of three main components, namely 1) a (virtual) prototype, 2) an oscilloscope, and 3) a host PC. This part discusses in detail the three components and how they interact with each other to deliver the proposed framework.
@@
The single-core, in-order, five-stage pipelined, 32-bit RISC-V CPU implements the base integer (I) and integer multiplication and division (M) RISC-V 32-bit ISA extensions [33]. Supporting only I and M extensions, i.e., the bare minimum extensions for an IoT-class CPU, provides the simplest and most observable setup, making it easier to carry out the side-channel analysis. A floating-point unit (FPU) can however be optionally instantiated as a functional unit of the CPU to expand the computing capabilities of the SoC [34, 35]. The interconnect consists of two separate 64-bit Wishbone data and instruction buses, both supporting single read and write transactions as well as burst ones, according to a modified Harvard architecture to minimize contention. The DFS actuator enables changing at run time the frequency of the clock signal fed to the CPU, whereas the rest of the SoC operates at a fixed clock frequency. A global debug unit and local debug units connected to the former in a point-to-point fashion compose the debug infrastructure of the SoC. The global debug unit interacts with the host PC through the system UART while two different channels support its communication with the rest of the SoC. Memory location of the memory-mapped bus slaves are accessed through Wishbone reads and writes, while dedicated lines connect the global debug unit to the two local ones that manage the CPU and DFS actuator, respectively.

### v2

@@ Abstract
Attacks based on side-channel analysis (SCA) pose a severe security threat to modern computing platforms, further exacerbated on IoT devices by their pervasiveness and handling of private and critical data. Designing SCA-resistant computing platforms requires a significant additional effort in the early stages of the IoT devices’ life cycle, which is severely constrained by strict time-to-market deadlines and tight budgets. This manuscript introduces a hardware-software framework meant for SCA research on FPGA targets. It delivers an IoT-class system-on-chip (SoC) that includes a RISC-V CPU, provides observability and controllability through an ad-hoc debug infrastructure to facilitate SCA attacks and evaluate the platform’s security, and streamlines the deployment of SCA countermeasures through dedicated hardware and software features such as a DFS actuator and FreeRTOS support. The open-source release of the framework includes the SoC, the scripts to configure the computing platform, compile a target application, and assess the SCA security, as well as a suite of state-of-the-art attacks and countermeasures. The goal is to foster its adoption and novel developments in the field, empowering designers and researchers to focus on studying SCA countermeasures and attacks while relying on a sound and stable hardware-software platform as the foundation for their research.
@@ I Introduction
Hundreds of billions of Internet of Things (IoT) devices are getting more and more able to autonomously make decisions thanks to artificial intelligence. As they get more capable, pervasive, and interconnected through 5G and 6G networks, IoT devices continuously collect, process, and exchange sensitive and critical data, making it paramount to consider security during their design and whole life cycle [1].
Relying on the theoretical security provided by traditional cryptography solutions such as AES, RSA, and SHA-3, and by secure communication protocols such as TLS and SSH is, however, insufficient to guarantee IoT device security. Side-channel analysis (SCA) attacks target information collected from executing a specific implementation of a cryptographic scheme or security protocol rather than flaws in their specification or the theoretical problems at their foundation [2]. Such so-called side-channel information includes factors like the time taken for cryptographic computations, power consumption, electromagnetic emissions, and noise generated during execution. While some fault-injection attacks disrupt the target device’s normal operation, many passive side-channel attacks, such as differential power analysis (DPA), remain undetectable by the system under attack.
While SCA attacks increasingly become serious security threats to IoT devices, and in particular to those that operate in public spaces and are accessible by anyone, the computing platforms that power these devices and that handle sensitive and critical data are, however, often not designed to protect against them. Commercially available microcontroller-based system-on-chips (SoCs), which provide computing and connectivity capabilities to IoT devices, often execute cryptographic applications despite not being designed with security as a primary focus, primarily due to cost-saving considerations [3]. Meanwhile, strict time-to-market pressures, combined with stringent constraints on energy, power consumption, and area, along with the rapidly evolving application domain, drive a shift towards flexible and general-purpose IoT platforms that may support operating systems to facilitate programmability. However, this trend towards general-purpose IoT devices results in more complex computing platforms, making side-channel vulnerability analysis increasingly challenging and necessitating the adoption of novel and efficient analysis frameworks.
In this context, the open literature lacks a comprehensive hardware-software solution to address the challenges of SCA security, particularly for modern RISC-V-based platforms, which are promising candidates for secure-aware designs due to their open and royalty-free instruction set architecture (ISA). The few RISC-V security-oriented solutions, e.g., OpenTitan [4], are not meant as general-purpose computing platforms and do not support the development of novel SCA attack and defense methodologies, while others focus their research effort solely on the hardware side while disregarding the software framework [5] or vice versa.
@@
This manuscript introduces JARVIS, an open-source research framework for SCA on FPGA-based IoT-class computing platforms, that encompasses a complete SoC and a software toolchain for SCA attacks and countermeasures, delivering a comprehensive hardware-software solution that is, to the best of our knowledge, currently lacking in the open literature.
The framework provides an IoT-class SoC that includes a CPU based on the RISC-V ISA, dedicated hardware to enable the implementation of state-of-the-art SCA countermeasures, an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, and support for the open source FreeRTOS real-time operating system (RTOS). A complete flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify SCA vulnerabilities and leakage sources.
The goal of empowering designers and researchers to focus solely on studying SCA countermeasures is achieved through three main contributions.
  1. 1.
    Capability to identify leakage sources. Hardware security requires not just identifying the presence of a vulnerability, but also pinpointing the source of such leakage. The computing platform has a minimal architecture to expose the least leakage sources and make them eventually simpler to identify, in addition to being easier to emulate and cheaper in terms of area and power cost, while a one-to-one match between the prototyped platform and its emulated counterpart is enforced through dedicated hardware mechanisms to enable the collection of the most accurate side-channel signal.
  2. 2.
    Usability for research purposes. The framework is released open source [^1] Sources available at https://github.com/hardware-fab/JARVIS., including hardware and software, from the SoC to the software scripts driving the SoC configuration, compilation, prototype execution and emulation, oscilloscope measurement, and state-of-the-art SCA attacks, to foster its adoption in research settings. Carrying out an analysis through the JARVIS framework not only provides the traditional power, performance, and area (PPA) quality metrics, but it extends them by adding the security dimension.
  3. 3.
    Complete SCA ecosystem. The framework includes dedicated hardware and software support for state-of-the-art SCA attacks and countermeasures that can be employed out of the box to directly compare with solutions from the literature as well as evaluate and enforce the SCA security of the computing platform. The experimental evaluation in this manuscript demonstrates the framework’s capabilities and showcases the SCA attacks and countermeasures included as part of its open-source release.
@@
The rest of this manuscript is organized into five parts. Section II provides a background on SCA attacks and countermeasures and an overview of the state-of-the-art RISC-V computing platforms and frameworks for SCA attacks and countermeasures. Section III describes the JARVIS framework and the hardware-software infrastructure that implements it. Section IV details the key hardware and software aspects that enable the SCA attack and countermeasure capabilities of the framework, Section V showcases the framework capabilities through a comprehensive experimental evaluation. Finally, Section VI draws conclusions and discusses the future works and developments on the proposed framework.
@@ II-A Side-channel analysis attacks and countermeasures
SCA attacks can be classified as either non-profiled or profiled. Non-profiled ones, such as those based on DPA [6] and correlation power analysis (CPA) [7], only target side-channel information that is obtained from the specific device under attack while it is performing a computation, and then attempt to recover the corresponding secret key by leveraging the statistical correlation between the measured side-channel signal and the data being processed, exploiting a partial knowledge of the latter. Profiled attacks [8] are carried out instead by initially making use of a replica of the device to be attacked to identify and fine-tune the side-channel leakage model and later employing such model to attack the actual target device, under the assumption that the latter and its replica share identical or at least similar leakage models [9]. Whereas SCA attacks have traditionally employed statistical techniques, machine- and deep-learning approaches recently emerged as promising research avenues for more capable attacks, ranging from identification of cryptographic operations from side-channel power traces [10, 11] to learning-based non-profiled [12] and profiled [13, 14] attacks.
@@
Figure 1: High-level flow of the JARVIS framework.
@@ II-B State-of-the-art RISC-V computing platforms
On the one hand, a variety of RISC-V-based microcontroller-class SoCs is available, such as PULPino [20] and GAP-8 [21], while ultra-low-power platforms also leverage RISC-V in multi-core heterogeneous architectures, e.g., Mr. Wolf [22] and HERO [23], and SoCs coupling hard-core CPU clusters with programmable FPGA logic, e.g., Microchip’s PolarFire SoC [24]. Such solutions focus on delivering good performance with low energy and power consumption but do not address security, especially SCA-related aspects. Moreover, due to their nature, designers can not modify ASIC cores to implement SCA countermeasures not originally included, drastically limiting their usability for comprehensive SCA research.
On the other hand, multiple frameworks such as ESP [25] and Chipyard [26] have emerged in the literature for the agile development of heterogeneous multi- and many-core SoCs that feature both general-purpose CPU cores and hardware accelerators [27]. However, they focus on the modularity and composability of the obtained computing platforms and on the fast prototyping of the latter on FPGA targets, while they do not consider instead any security aspects.
@@
Finally, the OpenTitan project [4] delivers a hardware root of trust (RoT) centered around a RISC-V core, not providing instead either a complete SoC platform or a comprehensive hardware-software framework for SCA-related research.
There also exists a variety of both RISC-V resource-constrained and application-class cores in the open literature, e.g., CORE-V CV32E40P [28] and CVA6 [29], UCB’s Rocket [30], and Frontgrade Gaisler NOEL-V [31], that are however of limited use, when considered on their own, in providing a full platform for SCA research. State-of-the-art architectural simulators such as gem5 [32] and RISC-V-specific GVSoC [33] operate instead at a high level of abstraction, drastically shrinking the time required to simulate a computing platform but conversely making them unsuitable to produce information that can be effectively exploited by SCA attacks.
II-C State-of-the-art SCA frameworks
The literature offers few solutions that are meant for the SCA of RISC-V-based computing platforms.
lowRISC provides an open-source SCA setup for OpenTitan [34] allows instantiating the OpenTitan RoT on a target NewAE board and assessing the resistance of its cryptographic accelerators to power SCA attacks. The OpenTitan platform occupies however a large number of FPGA resources and requires therefore targeting large and expensive FPGAs, e.g., AMD Kintex-7 ones, leading to non-negligible costs, and does not offer any actuators that are suitable for protection against SCA attacks. Moreover, the size and complexity of OpenTitan makes it complex to modify the given design with the goal of implementing novel SCA countermeasures, and the SCA setup [34] lacks the ability to easily identify the leakage source in the microarchitectural design.
ASIC-based solutions, that include mounting on NewAE’s CW308 board an FE310-G002 daughter board [35] which features a hard-core SiFive 32-bit RISC-V CPU but lacks any SCA protection, have two main drawbacks. On the one hand, pinpointing the side-channel leakage within the microarchitecture is not easily achievable due to the limited observability and closed-source nature of the computing platform. On the other hand, the researcher can not easily modify the hardware to implement SCA countermeasures.
Finally, NewAE’s CW305, CW310, and CW340 boards also allow the users to flash their own custom bitstreams, including RISC-V-based computing platforms. This requires however either porting a third-party computing platform to the new target or designing a new one, and both paths notably require a large amount of effort.
@@ III Framework
The JARVIS framework enables thoroughly evaluating the vulnerability to SCA attacks of cryptographic applications executed on an IoT-class SoC that implements countermeasures through dedicated hardware support. The possibility to accurately match the emulated and prototype execution of the computing platform allows not only detecting information leakage, but even more importantly identifying its sources, which is paramount to fixing the vulnerabilities of the SoC during the design phase. Section III-A outlines the high-level flow for the proposed framework, while Section III-B discusses how such flow is implemented in a hardware-software infrastructure.
@@ III-A High-level flow
The JARVIS framework, depicted in Figure 1, can be described as a sequence of three main phases devoted to 1) configuringthe hardware-software setup, 2) measuringSCA-related information from simulation and prototype execution, and 3) analyzingsuch information to detect leakage and identify its sources. Remarkably, the open-source nature of the framework, as well as the adoption of standard languages for all the inputs and its internals, enable the users to tinker, adapt, and tailor it to their needs. Software scripts automate the execution of each of the three phases, which are discussed in more detail in the rest of this section.
@@ III-A1 Configure
The high-level flow starts with the configuration of the hardware and software parts of the target to be simulated and prototyped. The SoC is thus configured by the user and is implemented according to the standard hardware design flow, while the applications and, optionally, the RTOS are compiled to produce an executable to be run by the CPU.
@@
During this phase, the source code for the applications provided by the user of the JARVIS framework is compiled, through a compiler toolchain for RISC-V, to be executed on the target platform, both in simulation and on the prototype. The compilation produces an executable file for the compiled applications and, optionally, the RTOS, meant to be executed by the computing platform both in simulation and on the prototype FPGA. The boundaries for the time window of interest in the execution of the application, e.g., the computation of a specific cryptographic kernel of which to evaluate SCA resistance, are also extracted from the executable file.
@@ III-A3 Analyze
The final phase of the JARVIS framework foresees the analysis of the measurements from the previous phase, i.e., the switching activity of the simulation and the power consumption trace of the board execution, both corresponding to the time window of interest and synchronized with each other. Two degrees of analysis are supported by the synchronization and alignment between the switching activity and the power trace. On the one hand, SCA leakage and the corresponding time instant are detected from the physical power traces. On the other hand, and more importantly, such alignment helps to identify the specific signals of the design under analysis which are responsible for the SCA leakage, providing an in-depth understanding of the leakage behavior that drastically eases the subsequent ad-hoc design of effective countermeasures. A set of state-of-the-art SCA techniques, ranging from CPA and template to ML-based ones, is therefore provided to evaluate the SCA security of the platform and the application executed on top of it. Remarkably, the analysis phase also produces the traditional power, performance, and area (PPA) metrics.
Figure 2: Hardware-software infrastructure that implements the JARVIS framework.
@@ III-B Hardware-software infrastructure
The hardware-software infrastructure depicted in Figure 2 realizes the high-level flow described in Section III-A by means of three main components, namely 1) a (virtual) prototype, 2) an oscilloscope, and 3) a host PC. This part discusses in detail the three components and how they interact with each other to deliver the JARVIS framework.
@@
The single-core, in-order, five-stage pipelined, 32-bit RISC-V CPU implements the base integer (I) and integer multiplication and division (M) RISC-V 32-bit ISA extensions [38]. Supporting only I and M extensions, i.e., the bare minimum extensions for an IoT-class CPU, provides the simplest and most observable setup, making it easier to carry out the side-channel analysis. A floating-point unit (FPU) [39, 40, 41] can be optionally instantiated as a functional unit of the CPU to expand the computational capabilities of the SoC, enabling the execution of floating-point-intensive applications without affecting the overall side-channel resistance of the computing platform. The interconnect consists of two separate 64-bit Wishbone data and instruction buses, both supporting single read and write transactions as well as burst ones, according to a modified Harvard architecture to minimize contention. The DFS actuator enables changing at run time the frequency of the clock signal fed to the CPU, whereas the rest of the SoC operates at a fixed clock frequency. A global debug unit and local debug units connected to the former in a point-to-point fashion compose the debug infrastructure of the SoC. The global debug unit interacts with the host PC through the system UART while two different channels support its communication with the rest of the SoC. Memory location of the memory-mapped bus slaves are accessed through Wishbone reads and writes, while dedicated lines connect the global debug unit to the two local ones that manage the CPU and DFS actuator, respectively.